– General Linux 2 – Fundamentals of TCP/IP [4]

(Linux Professional Institute Certification)

a

   .~.
   /V\      by: geoffrey robertson
  // \\         geoffrey@zip.com.au
  @._.@
$Id: gl2.112.1.slides.tex,v 1.1 2003/09/10 01:40:47 geoffr Exp $

aCopyright \begin{matrix}c \\ ◯ \end{matrix} 2003 Geoffrey Robertson. Permission is granted to make and distribute verbatim copies or modified versions of this document provided that this copyright notice and this permission notice are preserved on all copies under the terms of the GNU General Public License as published by the Free Software Foundation—either version 2 of the License or (at your option) any later version.

(2.2) Networking Fundamentals [14]

1.112.1

Fundamentals of TCP/IP [4] 1.112.3 TCP/IP configuration and troubleshooting [7] 1.112.4 Configure Linux as a PPP client [3]

Fundamentals of TCP/IP [4]

Objective

Candidates should demonstrate a proper understanding of network fundamentals. This objective includes the understanding of IP-addresses, network masks and what they mean (i.e. determine a network and broadcast address for a host based on its subnet mask in ”dotted quad” or abbreviated notation or determine the network address, broadcast address and netmask when given an IP-address and number of bits). It also covers the understanding of the network classes and classless subnets (CIDR) and the reserved addresses for private network use. It includes the understanding of the function and application of a default route.

It also includes the understanding of basic internet protocols (IP, ICMP, TCP, UDP) and the more common TCP and UDP ports (20, 21, 23, 25, 53, 80, 110, 119, 139, 143, 161).

Key files, terms, and utilities

/etc/services
ftp
telnet
host
ping
dig
traceroute
whois

Resources of interest

TCP/IP Illustrated, Volume 1 The Protocols

by W. Richard Stevens

Addison Wesley

LPIC 1 Certification Bible Angie Nash and Jason Nash Hungry Minds

IP addressing

This objective includes the understanding of IP-addresses, network masks and what they mean (i.e. determine a network and broadcast address for a host based on its subnet mask in ”dotted quad” or abbreviated notation or determine the network address, broadcast address and netmask when given an IP-address and number of bits).

IP Address Classes (Classic)

Class A - 255.0.0.0

\underline{0}0000000.00000000.00000000.00000000 - 0.0.0.0
\underline{0}1111111.11111111.11111111.11111111 - 127.255.255.255

Class B - 255.255.0.0

\underline{10}000000.00000000.00000000.00000000 - 128.0.0.0
\underline{10}111111.11111111.11111111.11111111 - 191.255.255.255

Class C - 255.255.255.0

\underline{110}00000.00000000.00000000.00000000 - 192.0.0.0
\underline{110}11111.11111111.11111111.11111111 - 223.255.255.255

IP Address Classes (Classic)

Class A - 255.0.0.0

\underline{0}0000000.00000000.00000000.00000000 - 0.0.0.0
\underline{0}1111111.11111111.11111111.11111111 - 127.255.255.255

Class B - 255.255.0.0

\underline{10}000000.00000000.00000000.00000000 - 128.0.0.0
\underline{10}111111.11111111.11111111.11111111 - 191.255.255.255

Class C - 255.255.255.0

\underline{110}00000.00000000.00000000.00000000 - 192.0.0.0
\underline{110}11111.11111111.11111111.11111111 - 223.255.255.255

8-a

IP Address Classes (Classic)

Class A - 255.0.0.0

\underline{0}0000000.00000000.00000000.00000000 - 0.0.0.0
\underline{0}1111111.11111111.11111111.11111111 - 127.255.255.255

Class B - 255.255.0.0

\underline{10}000000.00000000.00000000.00000000 - 128.0.0.0
\underline{10}111111.11111111.11111111.11111111 - 191.255.255.255

Class C - 255.255.255.0

\underline{110}00000.00000000.00000000.00000000 - 192.0.0.0
\underline{110}11111.11111111.11111111.11111111 - 223.255.255.255

8-b

IP Address Classes (Classic)

Class A - 255.0.0.0

\underline{0}0000000.00000000.00000000.00000000 - 0.0.0.0
\underline{0}1111111.11111111.11111111.11111111 - 127.255.255.255

Class B - 255.255.0.0

\underline{10}000000.00000000.00000000.00000000 - 128.0.0.0
\underline{10}111111.11111111.11111111.11111111 - 191.255.255.255

Class C - 255.255.255.0

\underline{110}00000.00000000.00000000.00000000 - 192.0.0.0
\underline{110}11111.11111111.11111111.11111111 - 223.255.255.255

8-c

IP Address - Loopback

Reserved Space 127.0.0.0 — 127.255.255.255 127.0.0.1 localhost

IP Address - Private Networks

There are IP ranges set aside for privite address spaces. These should not be made visible on the internet.

Class A

10.0.0.0 --- 10.255.255.255

Class B

172.16.0.0 --- 172.32.255.255

Class C

192.168.0.0 --- 192.168.255.255

IP Address - Subneting

Network: 192.168.192.0
Subnet:  255.255.255.224

IP Address - Default Route

# route add default gw 192.168.1.1

DoD Layer Model

Application ftp, telnet, mail, http protocols Transport TCP , UDP protocols Network IP, ICMP, IGMP protocols Link Ethernet, Token Ring, FDDI

Basic Internet Protocols

IP ICMP TCP UDP

Ports and Port Numbers

Listing the Ports

$ less /etc/services
ftp               21/tcp
ftp               21/udp           fsp fspd
ssh               22/tcp           # SSH Remote Login Protocol
ssh               22/udp           # SSH Remote Login Protocol
telnet           23/tcp
telnet           23/udp
# 24 - private mail system
smtp              25/tcp           mail
smtp              25/udp           mail
time              37/tcp           timserver

Ports and Port Numbers

FTP 20, 21 Telnet 23 SSH 22 smtp 25 DNS 53 http 80 pop3 110 nntp 119 netbios 137, 138, 139 imap2 143 snmp 161

Port Number vRanges

1-255

Original reserved ports (till 1992) (256-1023 UNIX)

1-1023

Well Known or Famous Port Numbers - Reserved

1024-65535

Unprivileged

The End √ • (2.2) Networking Fundamentals [14] . . . . . . . . 2 √ • Fundamentals of TCP/IP [4] . . . . . . . . . . . . . 3 √ • IP addressing . . . . . . . . . . . . . . . . . . . . . 7 √ • IP Address Classes (Classic) . . . . . . . . . . . . . 8 √ • IP Address - Loopback . . . . . . . . . . . . . . . . 9 √ • IP Address - Private Networks . . . . . . . . . . . . 10 √ • IP Address - Subneting . . . . . . . . . . . . . . . . 11 √ • IP Address - Default Route . . . . . . . . . . . . . . 12 √ • DoD Layer Model . . . . . . . . . . . . . . . . . . 13 √ • Basic Internet Protocols . . . . . . . . . . . . . . . 14 √ • Ports and Port Numbers . . . . . . . . . . . . . . . 15 √ • Ports and Port Numbers . . . . . . . . . . . . . . . 16 √ • Port Number vRanges . . . . . . . . . . . . . . . . 17